INVARIANT STEIN DOMAINS IN STEIN SYMMETRIC SPACES AND A NON-LINEAR COMPLEX CONVEXITY THEOREM

SIMON GINDIKIN BERNHARD KRÖTZ

Abstract. We prove a complex version of Kostant’s non-linear convexity theorem. Applications to the construction of G-invariant Grauert tubes of Riemannian symmetric G/K spaces are given.